3CH4(g)→C3H8(g)+2H2(g)

Calculate ΔG at 298 K if the reaction mixture consists of 41 atm of CH4, 0.013 atm of C3H8, and 2.0×10−2 atm of H2.
Answer #1

given

3CH4(g)→C3H8(g) + 2H2(g)

PC3H8= 0.013 atm

PH2 = 2.0×10−2 atm

PCH4 = 41 atm

we know that

ΔG = ΔG0 + RTlnQ

lets calculate ΔG0of the reaction

ΔG0rxn = ΔG0 f(products) - ΔG0 f(reactants)

= (ΔG0 f(C3H8(g)) + 2*ΔG0 f(H2(g)) ) – (3*ΔG0 f(CH4(g)))

Now

ΔG0 f(C3H8(g)) = -24.40 kJ/mol

ΔG0 f(CH4(g)) = -50.80 kJ/mol

ΔG0 f(H2(g)) = 0 kJ/mol

ΔG0rxn = (-24.40 kJ/mol – 2*0) – (3*-50.80 kJ/mol)

= -24.40 kJ/mol + 152.40 kJ/mol

= 128 kJ/mol

So

ΔG0 = 128 kJ/mol

Q = Pproducts/Preactants

= (PC3H8) * (PH2)2 / (PCH4)3

= (0.013)* (2.0×10−2)2 / 413

= 0.013 * 4*10-4 / 68921

= 7.5 * 10-11

ln(Q) = -23.3

now

ΔG = ΔG0 + RTlnQ

= 128 kJ/mol + (8.314 J/mol K * 298 K * -23.3)

= 128 kJ/mol – 57.727 kJ/mol

= 70 kJ/mol